英語 (English)

主題別英語 (Thematic English) (総科 ((人文・社創・総理 B)1 年))

ラックストンロバート · 非常勤講師 1 単位後期水 1 ・ 2

(平成 19 年度以前の授業科目: 『主題別英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (2)』 )

【授業の目的】This class will take a thematic approach to the learning of conversational English. Themes include: excursions, shopping, dining, and personal history.

【授業の概要】 Pronunciation practice, thematic lessons, role-plays and short skits. 【キーワード】 [キーワード] 【先行科目】 [先行科目] 【関連科目】 [関連科目] 【到達目標】 Students will improve there English by considering different situations

and themes. 【授業の計画】

1. Introductions

2. Unit 6.1

3. Unit 6.2

4. Unit 7.1

5. Unit 7.2

6. Unit 8.1

7. Quiz One and Skit Preparation

8. Skits

9. Unit 9.1

10. Unit 9.2

11. Unit 10.1

12. Unit 10.2

13. Unit 11.1

14. Unit 11.2

15. Quiz Two and Skit Preparation

16. Skit Day 【教科書】 Jack C. Richards: New Person to Person, Student Book 1 【参考書等】 [参考資料] 【成績評価の方法】 Quiz 1 25%; Skit 1 25%; Quiz 2 25%; Skit 1 25% 【再試験の有無】 [再評価] 【受講者へのメッセージ】 [注意]

【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=221152 【連絡先 (オフィスアワー・研究室・ E メールアドレス)】 ⇒ ラックストン .

English

Thematic English (総科 ((人文・社創・総理 B)1 年))

Robert Luxton · Part-time Lecturer 1 unit 後期水 1 ・ 2

(平成 19 年度以前の授業科目: 『主題別英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (2)』 )

Target⟩ This class will take a thematic approach to the learning of conversational

English. Themes include: excursions, shopping, dining, and personal history. Outline⟩ Pronunciation practice, thematic lessons, role-plays and short skits. �������⟩ [キーワード] Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ [関連科目] Goal⟩ Students will improve there English by considering different situations and

themes.

Schedule⟩

1. Introductions

2. Unit 6.1

3. Unit 6.2

4. Unit 7.1

5. Unit 7.2

6. Unit 8.1

7. Quiz One and Skit Preparation

8. Skits

9. Unit 9.1

10. Unit 9.2

11. Unit 10.1

12. Unit 10.2

13. Unit 11.1

14. Unit 11.2

15. Quiz Two and Skit Preparation

16. Skit Day Textbook⟩ Jack C. Richards: New Person to Person, Student Book 1 Reference⟩ [参考資料] Evaluation Criteria⟩ Quiz 1 25%; Skit 1 25%; Quiz 2 25%; Skit 1 25% Re-evaluation⟩ [再評価] Message⟩ [注意]

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=221152 Contact (Office-Hour, Room, E-mail)⟩

⇒ Luxton .